# An additive two-level parallel variant of the DMRG algorithm with coarse-space correction

### v1

@@ 4. Numerical Experiments
A well-known property of the classical one-site DMRG algorithm is that if the ranks of the ansatz tensor-train decomposition are exactly the separation ranks of the sought-after minimizing tensor, then the DMRG algorithm converges in a single sweep [36]. A similar property also seems to hold for the two-site DMRG algorithm (although we are unaware of a formal proof). An interesting question to ask is if the A2DMRG Algorithm 2 satisfies a similar property. In order to answer this question, we show in Figure 8 the error plots for the classical DMRG and A2DMRG algorithms applied to the $H_{6}$ and $H_{8}$ molecule with no rank truncation. These plots indicate that as soon as the TT ranks of the ansatz are sufficiently close to the maximal rank $\mathbb{r}=64$ for $H_{6}$ and $\mathbb{r}=256$ for $H_{8}$, the A2DMRG algorithm converges to within machine precision in a single global iteration. We remark that this property does not seem to be satisfied by tensor minimization algorithms based on Riemannian optimization.
Figure 8. Numerical evidence for the convergence of the classical DMRG and A2DMRG algorithms in a single sweep when the tensor train ranks of the ansatz equal the separation ranks of the sought-after minimizer. The tolerances for the eigenvalue solvers were set to $10^{-10}$, and the algorithms were run until the relative energy difference between successive sweeps/global iterations was smaller than $10^{-10}$. The molecular system here are $H_{6}$ (above) for which $d=12$ and $H_{8}$ (below) for which $d=16$.
@@ 5. Discussion and Perspective
In light of our numerical experiments and theoretical considerations, the additive two‐level DMRG algorithm (A2DMRG) opens several promising avenues for both analysis and application. First, while our coarse‐space correction clearly accelerates convergence in practice, a rigorous study of the convergence properties, and in particular how the choice of coarse subspace impacts convergence rates on strongly correlated systems, remains to be undertaken. Second, the intrinsic parallelism of local solves suggests that multilevel or adaptive Schwarz strategies (where coarse spaces are dynamically enriched based on algebraic error indicators) could yield further speedups and reliability. Addressing these points will deepen our understanding of two-level additive Schwarz methods in tensor networks.
@@ Appendix A Computational Scaling of the A2DMRG Algorithm
The aim of this section is to discuss in more detail the implementation and computational scaling of various steps in the additive two-level DMRG Algorithm 2 when the underlying energy functional $\mathcal{J}\colon U\subset\mathbb{R}^{n_{1}\times\ldots\times n_{d}}\rightarrow\mathbb{R}$ takes the form of a Rayleigh quotient as given by Equation (2), i.e.,
$\forall\mathbb{U}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\setminus\{0\}\colon\qquad\mathcal{J}(\mathbb{U})=\frac{\langle\mathbb{U},\mathbb{A}\mathbb{U}\rangle}{\|\mathbb{U}\|^{2}},$
@@
As is standard in the literature, we assume that the tensor operator $\mathbb{A}\colon\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ $\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ possesses a tensor train decomposition. In other words, viewing $\mathbb{A}$ as an order $d$-tensor in the space $\mathbb{R}^{(n_{1}\times n_{1})\times\ldots\times(n_{d}\times n_{d})}$, we assume that there exists some vector of ranks $\mathbb{R}=(\mathbb{R}_{0},\ldots\mathbb{R}_{d})\in\mathbb{N}^{d+1}$ with $\mathbb{R}_{0}=\mathbb{R}_{d}=1$ and $\{\mathbb{A}_{j}\}_{j=1}^{d},~{}\mathbb{A}_{j}\in\mathbb{R}^{R_{j-1}\times n_{j}\times n_{j}\times R_{j}}$ such that
@@
::=:= $\tau(\mathbb{A}_{1},\ldots,\mathbb{A}_{d})\big{(}(x_{1},x_{1}^{\prime}),\ldots,(x_{d},x_{d}^{\prime})\big{)}$
::=:= $\mathbb{A}_{1}(:,x_{1},x_{1}^{\prime},:)\mathbb{A}_{2}(:,x_{2},x_{2}^{\prime},:)\ldots\mathbb{A}_{d}(:,x_{d},x_{d}^{\prime},:)$
@@ A.1. Computational Scaling of Classical DMRG Algorithms
Let us first recall the implementation of the classical one-site DMRG  Algorithm 1. We assume that we have at hand a $j$-orthogonal tensor train decomposition $\mathbb{X}=(\mathbb{X}_{1},\ldots,\mathbb{X}_{d})$ for $j\in\{2,\ldots,d-1\}$ (the cases $j=1,d$ being simpler), and we are interested in performing the $j^{\rm th}$ one-site micro-iteration in the course of a left-to-right sweep. To do so we are required to perform matrix-vector products of the form
@@
Here, $\mathbb{A}_{j}^{\rm 1site,mat}\in\mathbb{R}^{r_{j-1}n_{j}r_{j}\times r_{j-1}n_{j}r_{j}}$ is the matricization of the operator $\mathbb{P}_{\mathbb{X},j,1}^{*}\mathbb{A}\mathbb{P}_{\mathbb{X},j,1}$ introduced in Lemma 10, and we remind the reader that $\mathbb{P}_{\mathbb{X},j,1},~{}\mathbb{X}\in\mathcal{U}_{\mathbb{r}}$ is the retraction operator defined through Definition 7.
@@
$\sum_{\ell_{j}=1}^{n_{j}}\sum_{\ell_{j}^{\prime}=1}^{n_{j}}\mathbb{X}_{j}(k_{j-1},\ell_{j},k_{j})\mathbb{A}_{j}(K_{j-1},\ell_{j},\ell_{j}^{\prime},K_{j})\mathbb{X}_{j}(k_{j-1}^{\prime},\ell_{j}^{\prime},k_{j}^{\prime}),$
@@
Since $r_{0}=R_{0}=r_{d}=R_{d}=1$, we can identify $\mathbb{G}_{1}$ and $\mathbb{G}_{d}$ as elements of $\mathbb{R}^{r_{j}\times R_{j}\times r_{j}}$ and $\mathbb{R}^{r_{j-1}\times R_{j-1}\times r_{j-1}}$ respectively. With this construction in mind, we now define for each $j\in\{1,\ldots,d\}$, the $j^{\rm th}$ left environment $\mathbb{H}^{\rm L}_{j}\in\mathbb{R}^{r_{j}\times R_{j}\times r_{j}}$ as
@@
Here, the indices $k_{j},k_{j}^{\prime}\in\{1,\ldots,r_{j}\}$ and $K_{j}\in\{1,\ldots,R_{j}\}$. Similarly, we define for each $j\in\{1,\ldots,d\}$, the $j^{\rm th}$ right environment $\mathbb{H}^{\rm R}_{j}\in\mathbb{R}^{r_{j-1}\times R_{j-1}\times r_{j-1}}$ as
@@
where the indices $k_{j-1},k_{j-1}^{\prime}\in\{1,\ldots,r_{j_{1}}\}$ and $K_{j-1}\in\{1,\ldots,R_{j-1}\}$.
It can now readily be seen that the matrix-vector product (24) can be computed by contracting the left environment $\mathbb{H}^{\rm L}_{j-1}$ with $\mathbb{A}_{j},\mathbb{v}_{j}$ and the right environment $\mathbb{H}^{\rm R}_{j+1}$ for micro-iteration $j>1$ and by contracting $\mathbb{A}_{j},\mathbb{v}_{j}$ with the right environment $\mathbb{H}^{\rm R}_{j+1}$ for micro-iteration $j=1$. Moreover, the computational cost of this step, i.e., the cost of constructing the environments $H_{j-1}^{\rm L},H_{j+1}^{\rm R}$ and performing the necessary tensor contractions scales as
@@
Here, we denote $n=\max\{n_{j}\}_{j=1}^{d}$, $r=\max\{r_{j}\}_{j=1}^{d}$ and $R=\max\{R_{j}\}_{j=1}^{d}$, and we have assumed (as is typically the case in practice) that $r>nR$. We also emphasize that these costs are calculated under the assumption that the left environment $\mathbb{H}_{j-1}^{\rm L}$ is constructed by successively contracting the tensors $\{\mathbb{G}_{\ell}\}_{\ell=1}^{j-1}$ from left-to-right while the right environment $\mathbb{H}_{j+1}^{\rm R}$ is constructed by successively contracting the tensors $\{\mathbb{G}_{\ell}\}_{\ell=d}^{j+1}$ from right-to-left.
It follows that the dominant cost in computing the matrix vector product (24) is the construction of the left and right environments $\mathbb{H}^{\rm L}_{j-1}$ and $\mathbb{H}^{\rm L}_{j+1}$ respectively, which raises the question of whether these objects can be precomputed at the beginning of the sweep and updated on the fly after each micro-iteration. Recalling the one-site DMRG algorithm 1 and considering Equations (25)-(26) that define these environments, we see that the answer is yes in the following sense: In the course of a left-to-right sweep, as we transition from micro-iteration $j$ to $j+1$, the left environment $H_{j}^{\rm L}$ can be constructed from $H_{j-1}^{\rm L}$ simply by contracting with $\mathbb{A}_{j}$ and the newly obtained solution $\widetilde{\mathbb{X}}_{j}$ to the $j^{\rm th}$ one-site DMRG micro-iteration. Consequently, in the course of a left-to-right sweep, at any one-site DMRG micro-iteration $j\in\{1,\ldots,d\}$, the construction from scratch of the left-environment $\mathbb{H}_{j-1}$ can be replaced with a simple update step of reduced cost. This leads to the following computational scaling of the matrix vector product (24):
@@
Remaining within the framework of a left-to-right sweep as before, we see that the situation with the right environment is different since transitioning from micro-iteration $j$ to $j+1$ requires us to compute $\mathbb{H}^{\rm R}_{j+2}$, which, in view of Equation (26) cannot be obtained by updating $\mathbb{H}^{\rm R}_{j+1}$. On the other hand, it is also clear from the one-site DMRG algorithm 1 that the right environments $\{\mathbb{H}_{j}\}_{j>1}$ do not depend on the solutions $\{\widetilde{\mathbb{X}}_{j}\}_{\ell<j}$ obtained from the one-site DMRG micro-iterations in the current sweep. With this observation, we see that we have essentially two possibilities:
  • If the maximal rank parameters $r,R$ and tensor order $d$ are sufficiently small and the storage capacity is sufficiently large, then we can precompute all $d$ right-environments $\{\mathbb{H}^{\rm R}_{j}\}_{j>1}$ at the beginning of the left-to-right sweep. Similar to the updates of the left environments discussed above, the right environment $H_{j}^{\rm R}$ can be constructed by updating $H_{j-1}^{\rm R}$ at a cost of $\mathcal{O}(nr^{3}R)$. Consequently, the total computational cost of this pre-computation step is $\mathcal{O}(dnr^{3}R)$. The matrix-vector product $\mathbb{A}_{1}^{\rm 1site,mat}\mathbb{v}_{1}$ for the first micro-iteration can now be performed simply by contracting $\mathbb{A}_{1}$ and $\mathbb{v}_{1}$ with the right environment $\mathbb{H}^{\rm R}_{2}$. Once the first micro-iteration is performed, we define the first left-environment $\mathbb{H}_{1}^{\rm L}$ and then adaptively update $\mathbb{H}_{j}^{\rm L},~{}j>1$ in the course of the one-site DMRG sweep. Since, we can additionally store the left-environments $\{\mathbb{H}_{j}^{\rm L}\}_{j=1}^{d}$ computed in this manner in memory, and since the role of the left and right environments is inverted when we transition from a left-to-right to a right-to-left sweep (see Algorithm 1), the pre-computation of the right environments $\{\mathbb{H}^{\rm R}_{j}\}_{j>1}$ has to be done only once at the start of the one-site DMRG algorithm. Consequently, we obtain the following improvement to the computational cost (29) of a matrix-vector product in each one-site DMRG micro-iteration:
@@
    In this framework therefore, if we denote by $K$ the maximal number of matrix-vector products needed in all one-site DMRG micro-iteration and by $M$ the number of one-site DMRG sweeps, we obtain the following computational scaling of the one-site DMRG algorithm:
@@
  • Turning now to the case when the parameters $r,R,$ and $d$ are very large and the storage capacity is limited, we see that the $j^{\rm th}$ micro-iteration in a left-to-right sweep requires an on-the-fly computation of the right-environment $\mathbb{H}^{\rm L}_{j+1}$ with cost ${\rm Cost}_{j,\rm Right\;Env}$. In this case, the computational cost (29) cannot be further improved, and if we denote again by $K$ the number of matrix-vector products needed in each one-site DMRG micro-iteration and by $M$ the number of one-site DMRG sweeps, we obtain the following computational scaling:
@@
Several remarks are now in order. First, note that in the above calculations, we have not explicitly taken into account the computational cost of performing the QR decomposition after each one-site DMRG micro-iteration. Since this cost scales as $\mathcal{O}(nr^{3})$ per micro-iteration, we see that it is dominated by the cost of the other steps in the one-site DMRG algorithm. Second, we emphasize that the same considerations (with some obvious modifications) apply in the case of the two-site DMRG algorithm. In particular, if all $d$ left/right environments are not being stored due to memory constraints, then we obtain the following computational scaling of the two-site DMRG algorithm:
(34) ${\rm Cost}^{\rm construct\;env}_{\rm 2site,compute}$ $=\mathcal{O}\big{(}dnr^{3}R(nK+d)\big{)}$
@@
We end this subsection with a short discussion on the computational costs of the one-site and two-site DMRG algorithms in a multi-processor setting. Several strategies to parallelize various steps in these algorithms (such as the tensor contractions) have been proposed in the physics literature (see, e.g., [5, 44, 39, 26, 29]). However, with the exception of the work of White and Stoudenmire [39], these parallelization strategies are all based on so-called pipeline parallelism in the sense that they focus on parallelism within the various steps comprising a micro-iteration. To our knowledge the work of White and Stoudenmire [39] is the only attempt at parallelism within the left/right sweeps of the by fundamentally altering the structure of the two-site DMRG algorithm.

### v2

@@ 4. Numerical Experiments
Figure 6. Numerical evidence for the convergence of the classical DMRG and A2DMRG algorithms in a single half-sweep when the tensor train ranks of the ansatz equal the separation ranks of the sought minimizer. The tolerances for the eigenvalue solvers were set to $10^{-10}$, and the algorithms were run until the relative energy difference between successive half-sweeps/global iterations was smaller than $10^{-10}$. The molecular systems here are H6 (above) for which $d=12$ and H8 (below) for which $d=16$.
A well-known property of the classical one-site DMRG algorithm is that if the ranks of the ansatz tensor-train decomposition are exactly the separation ranks of the minimizing tensor, then the DMRG algorithm converges in a single half-sweep [38]. A similar property also seems to holds for the two-site DMRG algorithm. An interesting question to ask is if the A2DMRG Algorithm 2 satisfies a similar property. In order to answer this question, we show in Figure 6 the error plots for the classical DMRG and A2DMRG algorithms applied to the H6 and H8 molecule with no rank truncation. These plots indicate that as soon as the TT ranks of the ansatz are sufficiently close to the maximal rank $\mathbb{r}=64$ for H6 and $\mathbb{r}=256$ for H8, the A2DMRG algorithm converges to within machine precision in a single global iteration. We remark that this property does not seem to be satisfied by tensor minimization algorithms based on Riemannian optimization.
@@ 5. Discussion and Perspective
The additive two-level DMRG algorithm raises interesting questions from the perspective of both analysis and applications. On the one hand, a rigorous convergence analysis of the algorithm, including a study of how the choice of coarse space impacts convergence rates, remains to be undertaken. Moreover, the parallelism of local solves suggests that adaptive Schwarz strategies wherein the coarse space is dynamically enriched (possibly based on error indicators) may yield further speedups and accuracy. Addressing these points can improve our understanding of Schwarz methods in the context of tensor networks and will be the subject of future work.
@@ Appendix A Computational Scaling of the A2DMRG Algorithm
The aim of this section is to discuss in more detail the implementation and computational scaling of various steps in the two-level A2DMRG Algorithm 2 when the underlying energy functional $\mathcal{J}\colon U\subset\mathbb{R}^{n_{1}\times\ldots\times n_{d}}\rightarrow\mathbb{R}$ takes the form of a Rayleigh quotient as given by Equation (2), i.e.,
$\forall\mathbb{X}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\setminus\{0\}\colon\qquad\mathcal{J}(\mathbb{X})=\frac{\langle\mathbb{X},\mathbb{A}\mathbb{X}\rangle}{\|\mathbb{X}\|^{2}},$
@@
As is standard in the literature, we assume that the tensor operator $\mathbb{A}\colon\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ $\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ possesses a tensor train decomposition. In other words, viewing $\mathbb{A}$ as an order $d$-tensor in the space $\mathbb{R}^{(n_{1}\times n_{1})\times\ldots\times(n_{d}\times n_{d})}$, we assume that there exists some vector of ranks $\mathbb{R}=(R_{0},\ldots R_{d})\in\mathbb{N}^{d+1}$ with ${R}_{0}={R}_{d}=1$ and $\{\mathbb{A}_{j}\}_{j=1}^{d},~\mathbb{A}_{j}\in\mathbb{R}^{R_{j-1}\times n_{j}\times n_{j}\times R_{j}}$ such that
@@
$:=$ $\tau(\mathbb{A}_{1},\ldots,\mathbb{A}_{d})\big((x_{1},x_{1}^{\prime}),\ldots,(x_{d},x_{d}^{\prime})\big)$
$:=$ $\mathbb{A}_{1}(:,x_{1},x_{1}^{\prime},:)\mathbb{A}_{2}(:,x_{2},x_{2}^{\prime},:)\ldots\mathbb{A}_{d}(:,x_{d},x_{d}^{\prime},:)$
@@ A.1. Computational Scaling of Classical DMRG Algorithms
Let us first recall the implementation of the classical one-site DMRG  Algorithm 1. We assume that we have at hand a $j$-orthogonal tensor train decomposition $\mathbb{U}=(\mathbb{U}_{1},\ldots,\mathbb{U}_{d})$ for $j\in\{2,\ldots,d-1\}$ (the cases $j=1,d$ being simpler), and we are interested in performing the $j^{\rm th}$ one-site micro-iteration in the course of a left-to-right half-sweep. To do so we are required to perform matrix-vector products of the form
@@
Here, $\mathbb{A}_{j}^{\rm 1site,mat}\in\mathbb{R}^{r_{j-1}n_{j}r_{j}\times r_{j-1}n_{j}r_{j}}$ is the matricization of the operator $\mathbb{P}_{\mathbb{U},j,1}^{*}\mathbb{A}\mathbb{P}_{\mathbb{U},j,1}$ introduced in Lemma 10, and we remind the reader that $\mathbb{P}_{\mathbb{U},j,1},~\mathbb{U}\in\mathcal{U}_{\mathbb{r}}$ is the retraction operator defined through Definition 7.
@@
$\sum_{\ell_{j}=1}^{n_{j}}\sum_{\ell_{j}^{\prime}=1}^{n_{j}}\mathbb{U}_{j}(k_{j-1},\ell_{j},k_{j})\mathbb{A}_{j}(K_{j-1},\ell_{j},\ell_{j}^{\prime},K_{j})\mathbb{U}_{j}(k_{j-1}^{\prime},\ell_{j}^{\prime},k_{j}^{\prime}),$
@@
Since $r_{0}=R_{0}=r_{d}=R_{d}=1$, we can identify $\mathbb{G}_{1}$ and $\mathbb{G}_{d}$ as elements of $\mathbb{R}^{r_{1}\times R_{1}\times r_{1}}$ and $\mathbb{R}^{r_{d-1}\times R_{d-1}\times r_{d-1}}$ respectively. With this construction in mind, we now define for each $j\in\{1,\ldots,d-1\}$, the $j^{\rm th}$ left environment $\mathbb{H}^{\rm L}_{j}\in\mathbb{R}^{r_{j}\times R_{j}\times r_{j}}$ as
@@
Here, the indices $k_{j},k_{j}^{\prime}\in\{1,\ldots,r_{j}\}$ and $K_{j}\in\{1,\ldots,R_{j}\}$. Similarly, we define for each $j\in\{2,\ldots,d\}$, the $j^{\rm th}$ right environment $\mathbb{H}^{\rm R}_{j}\in\mathbb{R}^{r_{j-1}\times R_{j-1}\times r_{j-1}}$ as
@@
where the indices $k_{j-1},k_{j-1}^{\prime}\in\{1,\ldots,r_{j-1}\}$ and $K_{j-1}\in\{1,\ldots,R_{j-1}\}$.
It can now readily be seen that for micro-iteration $j\in\{2,\ldots d-1\}$, the matrix-vector product (24) can be computed by contracting the left environment $\mathbb{H}^{\rm L}_{j-1}$ with $\mathbb{A}_{j},\mathbb{v}_{j}$ and the right environment $\mathbb{H}^{\rm R}_{j+1}$. Moreover, for micro-iteration $j=1$ (respectively, $j=d$) we simply need to contract $\mathbb{A}_{j}$ and $\mathbb{v}_{j}$ with the right environment $\mathbb{H}^{\rm R}_{2}$ (respectively, the left environment $\mathbb{H}^{\rm L}_{d-1}$). The computational cost of this step, i.e., the cost of constructing the environments $\mathbb{H}_{j-1}^{\rm L},\mathbb{H}_{j+1}^{\rm R}$ and performing the necessary tensor contractions scales as
@@
Here, we denote $n=\max\{n_{j}\}_{j=1}^{d}$, $r=\max\mathbb{r}$ and $R=\max\mathbb{R}$, and we have assumed (as is the case for state-of-the-art quantum chemistry simulations) that $r>nR$. We also emphasize that these costs are calculated under the assumption that the left environment $\mathbb{H}_{j-1}^{\rm L}$ is constructed by successively contracting the tensors $\{\mathbb{G}_{\ell}\}_{\ell=1}^{j-1}$ from left-to-right while the right environment $\mathbb{H}_{j+1}^{\rm R}$ is constructed by successively contracting the tensors $\{\mathbb{G}_{\ell}\}_{\ell=d}^{j+1}$ from right-to-left.
It follows that the dominant cost in computing the matrix vector product (24) is the construction of the left and right environments $\mathbb{H}^{\rm L}_{j-1}$ and $\mathbb{H}^{\rm R}_{j+1}$ respectively, which raises the question of whether these objects can be precomputed at the beginning of the half-sweep and updated on the fly after each micro-iteration. Recalling the one-site DMRG Algorithm 1 and considering Equations (25)-(26) that define these environments, we see that the answer is yes in the following sense: In the course of a left-to-right half-sweep, as we transition from micro-iteration $j$ to $j+1$, the left environment $H_{j}^{\rm L}$ can be constructed from $\mathbb{H}_{j-1}^{\rm L}$ simply by contracting with $\mathbb{A}_{j}$ and the newly obtained solution $\widetilde{\mathbb{U}}_{j}$ to the $j^{\rm th}$ one-site DMRG micro-iteration. Consequently, in the course of a left-to-right half-sweep, at any one-site DMRG micro-iteration $j\in\{1,\ldots,d\}$, the construction from scratch of the left-environment $\mathbb{H}^{\rm L}_{j-1}$ can be replaced with a simple update step of reduced cost. This leads to the following computational scaling of the matrix vector product (24):
@@
Remaining within the framework of a left-to-right half-sweep as before, we see that the situation with the right environment is different since transitioning from micro-iteration $j$ to $j+1$ requires us to compute $\mathbb{H}^{\rm R}_{j+2}$, which, in view of Equation (26) cannot be obtained by updating $\mathbb{H}^{\rm R}_{j+1}$. On the other hand, it is also clear from the description of the DMRG algorithm 1 that the right environments $\{\mathbb{H}^{\rm R}_{j}\}_{j>1}$ do not depend on the solutions $\{\widetilde{\mathbb{U}}_{\ell}\}_{\ell<j}$ obtained from the one-site DMRG micro-iterations in the current half-sweep. With this observation, we see that we have essentially two possibilities:
  • If the maximal rank parameters $r,R$ and tensor order $d$ are sufficiently small and the storage capacity is sufficiently large, then we can precompute all $d$ right-environments $\{\mathbb{H}^{\rm R}_{j}\}_{j>1}$ at the beginning of the left-to-right half-sweep. Similar to the updates of the left environments discussed above, the right environment $\mathbb{H}_{j}^{\rm R}$ can be constructed by updating $\mathbb{H}_{j-1}^{\rm R}$ at a cost of $\mathcal{O}(nr^{3}R)$. Consequently, the total computational cost of this pre-computation step is $\mathcal{O}(dnr^{3}R)$. The matrix-vector product $\mathbb{A}_{1}^{\rm 1site,mat}\mathbb{v}_{1}$ in the first micro-iteration can now be performed simply by contracting $\mathbb{A}_{1}$ and $\mathbb{v}_{1}$ with the right environment $\mathbb{H}^{\rm R}_{2}$. Once the first micro-iteration is performed, we define the first left-environment $\mathbb{H}_{1}^{\rm L}$ and then adaptively update $\mathbb{H}_{j}^{\rm L},~j>1$ in the course of the one-site DMRG half-sweep. Since, we can additionally store the left-environments $\{\mathbb{H}_{j}^{\rm L}\}_{j=1}^{d}$ computed in this manner in memory, and since the role of the left and right environments is inverted when we transition from a left-to-right to a right-to-left half-sweep (see Algorithm 1), the pre-computation of the right environments $\{\mathbb{H}^{\rm R}_{j}\}_{j>1}$ has to be done only once at the start of the one-site DMRG algorithm. Consequently, we obtain the following improvement to the computational cost (29) of a matrix-vector product in each one-site DMRG micro-iteration:
@@
    In this framework therefore, if we denote by $K$ the maximal number of matrix-vector products needed in all one-site DMRG micro-iteration and by $M$ the number of one-site DMRG half-sweeps, we obtain the following computational scaling of the one-site DMRG algorithm:
@@
  • Turning now to the case when the parameters $r,R,$ and $d$ are very large and the storage capacity is limited, we see that the $j^{\rm th}$ micro-iteration in a left-to-right half-sweep requires an on-the-fly computation of the right-environment $\mathbb{H}^{\rm L}_{j+1}$ with cost ${\rm Cost}_{j,\rm Right\;Env}$. In this case, the computational cost (29) cannot be further improved, and if we denote again by $K$ the number of matrix-vector products needed in each one-site DMRG micro-iteration and by $M$ the number of one-site DMRG half-sweeps, we obtain the following computational scaling:
@@
Several remarks are now in order. First, note that in the above calculations, we have not explicitly taken into account the computational cost of performing the QR decomposition after each one-site DMRG micro-iteration. Since this cost scales as $\mathcal{O}(nr^{3})$ per micro-iteration, it is dominated by the cost of the other steps in the one-site DMRG algorithm. Second, we emphasize that the same considerations (with some obvious modifications) apply in the case of the two-site DMRG algorithm. In particular, if all $d$ left/right environments are not being stored due to memory constraints, then we obtain the following computational scaling of the two-site DMRG algorithm:
(34) ${\rm Cost}^{\rm construct\;env}_{\rm 2site,compute}$ $=\mathcal{O}\big(dnr^{3}RM(nK+d)\big)$
@@
We end this subsection with a short discussion on the computational costs of the one-site and two-site DMRG algorithms in a multi-processor setting. Several strategies to parallelize various steps in these algorithms (such as the tensor contractions) have been proposed in the physics literature (see, e.g., [5, 46, 41, 28, 31]). However, with the exception of the work of White and Stoudenmire [41], these parallelization strategies are all based on so-called pipeline parallelism in the sense that they focus on parallelism within the various steps comprising a micro-iteration. To our knowledge the work of White and Stoudenmire [41] is the only attempt at parallelism within the left/right half-sweeps by fundamentally altering the structure of the two-site DMRG algorithm.